Solve the recurrence relation A(n) = 6A(n - 1) - 11A(n - 2) + 6A(n - 3) subject to initial values A(1) = 2, A(2) = 6, A(3) = 20.

1
Expert's answer
2022-02-25T12:59:40-0500

The characteristic equation of the recurrence relation is


"r^3 - 6r^2+ 11r-6 = 0"

"(r-1)(r-2)(r-3)=0"

Hence


"a_n=b_1+b_22^n+b_33^n"

"a_1=2, 2=b_1+2b_2+3b_3"

"a_2=6, 6=b_1+4b_2+9b_3"

"a_3=20, 20=b_1+8b_2+27b_3"



"b_1+2b_2+3b_3=2"

"2b_2+6b_3=4"

"6b_2+24b_3=18"



"b_1+2b_2+3b_3=2"

"b_2=2-3b_3"

"2-3b_3+4b_3=3"

"b_1=1"

"b_2=-1"

"b_3=1"

The required solution is


"a_n=1-2^n+3^n"